OpenStudy (anonymous):

find all the real cube roots of -0.000125

OpenStudy (anonymous):

A. 0.0025 and -0.0025 B. -0.05and 0.05 C. -0.05 D. 0.0025

OpenStudy (kc_kennylau):

Here is what will be in my mind: \[\sqrt[3]{-0.000125}=\sqrt[3]{-10^{-6}\times5^3}=-10^{-2}\times5=-0.05\] It may seem to be more complicated, but it's what's going on in my mind :)

OpenStudy (kc_kennylau):

https://www.youtube.com/watch?v=DKh16Th8x6o This video would help you in answering more questions of the same kind :D

OpenStudy (anonymous):

lol yeah it does

OpenStudy (anonymous):

okay thanks!

OpenStudy (kc_kennylau):

no problem :)